AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 295     Entry time: Sun Feb 27 04:54:46 2022
Author: David, Pierre-Michell 
Subject: Current status 

Ring current 8e7 particles/s

Merger, Tape server (WS3): 1664 kB/s writing ok

Leakage current (WS4): 3.26 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

Attachment 1: 2022-02-27_05-56-59exra.png  157 kB  | Hide | Hide all
2022-02-27_05-56-59exra.png
Attachment 2: 2022-02-27_05-56-31exto.png  166 kB  | Hide | Hide all
2022-02-27_05-56-31exto.png
Attachment 3: 2022-02-27_05-56-13xytot.png  175 kB  | Hide | Hide all
2022-02-27_05-56-13xytot.png
ELOG V3.1.4-unknown